(a) Based on your understanding of the paper entitled “The Existence of a Stable Money Multiplier...

  1. (a) Based on your understanding of the paper entitled “The Existence of a Stable Money Multiplier in the Small Open Economy of Kazakhstan”, briefly describe and explain the money multiplier and the channels through which a change in the monetary base affects the money supply in an economy. What is the implication of money multiplier for monetary policy?
  1. Assuming that all transactions are conducted through commercial banks in Saudi Arabia, explain how much an initial deposit of 100,000 Rials in a first bank creates the money supply if SAMA sets the legally required ratio reserve at 15%?
Answer #1

Ans a.)

Money multiplier is the magnitude of the change in the money supply when the monetary base changes by one unit. In other words it tells by how much money supply of an economy would change if the monetary base changes by a certain proportion.

Monetary base can change through the following channels:

Open market operations:

Open market operations refers to the Purchase and sale of Government Bonds and securities by the central bank from general public and banks.

When the government buys the bonds from the banks, it gives some amount in return which leads to the increase in the monetary base. This increase in the monetary base further leads to the increase in the reserves of banks and it further increases their capacity to make loans which ultimately leads to the increase in the money supply in an economy. This is called the expansionary open market operations as the money supply is increasing.

Discount rate:

Account rate is the minimum rate that the fed charges from the other banks for the loan. When the discount rate is decreased it tends to increase the loans given to the commercial banks which further leads to increase in their Reserves. This increase in the reserves leads to the increase in the money supply in an economy.

Money Multiplier and the monetary policy:

We know that the money multiplier just multiplies the vase money and leass to the increase in the money supply in an economy. Money multiplier is a combination of various factors. For the sake of simplicity me only consider reserve ratio as of now.

Reserve ratio is the proportion of deposits that a bank is supposed to keep with it in the liquidity form.

Higher is the value of the money multiplier, higher would be the money supply in the economy and vice versa.

Ans b.) Reserve ratio = 15%

Money multiplier = 1/RR

Where , RR = reserve ratio

  Money multiplier = 1/(15/100) = 100/15 = 6.67

Change in money supply = changebase money × money multiplier

= 6.67 × 100000

Increase in money supply = $667,000